Carl invested $10,000, part in a CD paying 8.5% annual interest, the rest in a mutual fund paying 7% annual interest. The annual income from the CD is $200 more than the annual income from the mutual fund. How much money was put into each type of investment?
Answer by mananth(16946)

Part I 8.50% per annum
Part II 7.00% per annum

x+y=10000 ------------------------1
8.50%x-7.00%y=$200.00
Multiply by 100
8.5x-7y =$20,000.00--------2

Multiply (1) by -8.5
we get

-8.5x-8.5y=-85000.00

Add this to (2)

x-15.5y =-$65,000.00

divide by -15.5

y=$4,193.55 investment at7.00%
Balance $5,806.45 investment at 8.50%